Calculate the mass of a sample of CO2 with a volume of 0.575 at 50°C and 115 kPa


1
Expert's answer
2022-05-13T00:49:02-0400

PV=nRT

V=0.575 l=0.000575 m3

T=50°C=323 K

P=115000 Pa

R=8.31 J/(mol K)

n=PV/(RT)=115000x0.000575/(8.31x323)=0.02463 mole

M(CO2)=12+16x2=44 g/mole

m(CO2)=44x0.02463=1.084 g